OpenStudy (anonymous):

The sun’s mass is about 2.7 x 107 times greater than the moon’s mass. The sun is about 400 times farther from Earth than the moon. How does the gravitational force exerted on Earth by the sun compare with the gravitational force exerted on Earth by the moon?

OpenStudy (badhi):

for easiness take newtons gravitational equation as, \[F \propto \frac{M}{r^2}\] (since \(m\), \(G\) are common for both scenarios) for moon=> \[F_m\propto \frac{M_m}{r_m^2}\] for sun, \[F_s \propto \frac{M_s}{r_s^2} \] by dividing both, get the equation

OpenStudy (anonymous):

im a little lost

OpenStudy (badhi):

well this is a bit of a easy manipulation method, but if its confusing i'll describe, The gravitational field equation is,\[F=\frac{GmM}{r^2}\] If the gravity force between the moon and earth is \(F_m\) and mass of moon is \(M_m\) and the distance between moon and earth is \(r_m\) and the mass of the earth is \(m\) the above equation becomes,\[F_m = \frac{GmM_m}{r_m^2}= (Gm) \frac{M_m}{r_m^2}\] For force between the sun and earth is \(F_s\) and mass of the sun is \(M_s\) and distance is \(r_s\) , equation becomes,\[F_s = \frac{GmM_s}{r_s^2} = (Gm) \frac{M_s}{r_s^2}\] Is it clear so far?

OpenStudy (anonymous):

yes

OpenStudy (badhi):

now what you need is to compare \(F_s\) with \(F_m\) or get a value for \(\displaystyle \frac{F_s}{F_m}\) For that what should you do?

OpenStudy (anonymous):

divide F8/Fm

OpenStudy (badhi):

there you go.... any problems after that?

OpenStudy (anonymous):

only curious what do you plug into it so you can divide

OpenStudy (badhi):

first divide and see what terms get eliminated by division... Then try to plugin the values given as data... Remember what is given is as ratios too... For example, they say sun is \(2.7\times 10^7 \) times larger than moon in mass. Simply what that mean is, \(M_s = 2.7\times 10^7 M_m\)

OpenStudy (anonymous):

8/m?

OpenStudy (badhi):

whats that?

OpenStudy (anonymous):

F8/Fm F cancels rights

OpenStudy (badhi):

you mean \(\displaystyle \frac{F_s}{F_m} = \frac{s}{m}\) ??

OpenStudy (anonymous):

yah...is that wrong

OpenStudy (badhi):

oh dear... \(F_s\) and \(F_m\) are symbols as whole Just like \(A\) or \(B\) or any other symbols.. you jst cannot cut out parts of the symbol... simply what you are saying is like, \[\frac{V}{W} = \frac{1}{V}\] :P (by cutting out a apart of W with V since W is like two Vs)

OpenStudy (anonymous):

my bad so what do i do than

OpenStudy (badhi):

If you are not comfortable with the notation, Jst replace \(F_m,F_s, M_m, M_s, r_m, r_s\) with normal english letters such as \(a,b,c,d,e,f\) etc. and try out.. And please try to show what you've tried out

OpenStudy (anonymous):

no its not the symbols i just dont understand what im supposed to do to divide F8 by Fm

OpenStudy (badhi):

I have given you two equations for F_s and F_m. What you have to do is divide like, \(F_s = abc\), \(F_m = pqr\) \[\frac{F_s}{F_m} = \frac{abc}{pqr}\] Havent you done equation division before?

OpenStudy (anonymous):

idk

OpenStudy (badhi):

ok here is a problem solve it for me.. find \(\frac{F}{T}\) \(F=ba^2\) \(T=bc\)

OpenStudy (anonymous):

2b*c*a^2

OpenStudy (badhi):

I need the steps how you got there... not jst the answer..

OpenStudy (anonymous):

idk how to explain it i just factored it out

OpenStudy (anonymous):

i know if you divide a equation by an equation you actually combine and multiply the variables

OpenStudy (anonymous):

there are 2 b(s) hence 2b theres 1 c hence times c and 1 a^2 hence times a^2

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

so f8/fm becomes 2f8m

OpenStudy (anonymous):

\[2F _{8}m\]

OpenStudy (badhi):

well thats not how it works... I think you should work out more with your equations and symbols. before getting into this kinda applications.

OpenStudy (anonymous):

its the last question i have than im done :(

OpenStudy (badhi):

\[\frac{F}{T} = \frac{ba^2}{bc}=\frac{a^2}{c}\] when the left hand side is divided same goes to the right hand side.. then if there are common symbols they jst cross out jst like what happen to the b here

OpenStudy (badhi):

Sorry I cannot jst give out the answer to this...

OpenStudy (anonymous):

im not asking you to just help me understand what i need to do

OpenStudy (anonymous):

ohhh ok i get it

OpenStudy (badhi):

So would you mind showing us how to get the answer..

OpenStudy (anonymous):

yes but i just need to know what am i pluging in as the right hand side against F8/Fm

OpenStudy (anonymous):

nvm i know

OpenStudy (anonymous):

nvm i know what i need to plug in

OpenStudy (anonymous):

\[\frac{ F _{8} }{ Fm }=\frac{ \frac{ (Gm)M _{m} }{ r _{m}^2} }{ \frac{ (Gm)M _{8} }{ r _{8}^2 } }=\frac{ \frac{ M _{m} }{ r _{m}^2 } }{ \frac{ M _{8} }{ r _{8}^2 } }\]

OpenStudy (anonymous):

?

OpenStudy (badhi):

good

OpenStudy (badhi):

\[\frac{\left(\frac ab\right)}{\left( \frac c d\right)} = \frac a b \times\frac d c\] use this now

OpenStudy (anonymous):

\[\frac{ M _{m} }{ r _{m}^2 }*\frac{ M _{8} }{ r _{8}^2 }=2M _{m8}r _{m8}^4\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

?

OpenStudy (badhi):

note that when the division is changed to multiplication, c/d is changed to d/c...so what you've written there is wrong also... as i have mentioned before, \(M_s\times M_m\) cannot be further simplified to \(2M_{ms}\) (u can, maybe with some other algebra :P). It should jst be \(M_sM_m\) and treat \(M_s\) as a whole which cannot be separated and same goes to \(M_m\)

OpenStudy (anonymous):

ohh ok so how would I write it @Badhi

OpenStudy (badhi):

\[\frac{M_m}{r_m^2}\frac{r_s^2}{M_s}\] and then adjust it like this, \[\left(\frac{M_m}{M_s}\right)\left(\frac{r_s}{r_m}\right)^2\]